OpenStudy (anonymous):

The area of a rectangle is the product of width and length. If a rectangle has length 12m(sqrt21m) and width (sqrt16m) , which of the following is the area of the rectangle?

OpenStudy (anonymous):

Area of rectangle= Width x Length/ Length x Width (it's interchangeable and the answer will still remain the same. e.g. 3 x 4=12 and 4 x 3=12) \[12\sqrt{21} \times \sqrt {16}\] If this is how the numbers look like, then all you have to do from here is solve the equation. Do you need help with that?

OpenStudy (anonymous):

yah i need help with that

OpenStudy (anonymous):

144

OpenStudy (anonymous):

and the square root of 16?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Yup! Now, your equation would become \[12\sqrt{21} \times 4\] now, what is 4 x 12?

OpenStudy (anonymous):

48

OpenStudy (anonymous):

Right! :) Now, you have... \[48\sqrt{21}\] as your answer. (If you want a decimal, then I'd use the 144. However, I do not think you need to give a decimal answer so I'm leaving it in radical form.)

OpenStudy (anonymous):

Do you kind of get me? :/

OpenStudy (anonymous):

yah so that is the answer there is still work to be done

OpenStudy (anonymous):

that should be the answer unless your question asks to give the answer as a decimal.

OpenStudy (anonymous):

no decimal Thank You So Much!!! :)

OpenStudy (anonymous):

You're welcome! :) All the best!